# Two-particle calculations with quantics tensor trains - solving the parquet equations

### v2

@@ V.3 Parquet equation
The parquet equation (15) with its frequency shifts as in (16) can be solved entirely in QTT by first converting all the vertex functions $\Lambda,\Phi$ to the required channel as described in the previous section, then performing their summation and subtraction as shown in Ref. [34, Sec. 4.7]. The resultant QTTs for $F$ (and subsequently $\Gamma$ obtained from (13)) are then compressed to a maximum bond dimension $D_{\mathrm{max}}$ in $\mathcal{O}(D_{\mathrm{max}}^{3}R)$ computation time. Hence, the parquet equation has the same $\mathcal{O}(D_{\mathrm{max}}^{3}R)$ computational cost as channel transformations. Further investigation of error and runtime scaling can be found in App. D.
@@ V.4 Bethe–Salpeter equation
where $|_{\omega=\omega^{\prime\prime}}$ denotes the restriction of the result to $\omega=\omega^{\prime\prime}$. Note that Eq. (24a) has the structure of a matrix multiplication in the combined index $(\nu^{\prime},\omega^{\prime})$. Thus, this equation can be evaluated in QTT format through standard MPO-MPO contraction, as is illustrated in Fig. 8(b). Introducing dummy indices has a runtime and memory cost of $\mathcal{O}(D_{\mathrm{max}}^{2}R)$, which is much smaller than the cost of other steps in the algorithm.
@@ VI.1 Hubbard atom
A significant portion of the physics of strongly correlated electrons in the strong-coupling limit of the Hubbard model can already be captured by studying its atomic limit [1, 44]. This limit offers considerable simplification, as the vertex functions are analytically known and become independent of momentum [37]. Therefore, the Hubbard atom serves as an ideal first test case for exploring two-particle properties in strongly correlated electron systems using the QTT framework.
@@
Figure 11: (a) Maximum absolute normalized error $||\Delta\Gamma_{d}||_{\infty}:=||\Gamma_{d,\text{iterative-parquet}}-\Gamma_{d,\text{exact}}||_{\infty}/||\Gamma_{d,\text{exact}}||_{\infty}$ of the iterative parquet scheme with QTCI for the irreducible vertex $\Gamma_{d}$ compared to the exact values for various grid sizes ($2^{3R}$ grid points) in the case of $U=\beta=1$, plotted (a) as a function of $D_{\mathrm{max}}$ after 30 iterations, and as a function of iteration number for (b) $D_{\mathrm{max}}=100$ and (c) $D_{\mathrm{max}}=200$. Dashed lines indicate the results from the dense grid calculation without QTCI. (d) The maximum absolute normalized error of $\Gamma_{d}$, shown close to the first divergence, for $\beta=1.55,U=2.3$ and $D_{\mathrm{max}}=100$.
@@
Next, we show that computations can also be performed for a more challenging case with the parameters $\beta=1.55,U=2.3$, which are chosen in the same way as in Ref. [13]. This case is interesting since $\beta U=3.565$ is very close to the point, where the irreducible vertex in the density channel diverges ($\beta U\approx 3.628$[48, 49, 50, 51, 52]. We show the results of these calculations in Fig. 11(d) for a maximum bond dimension $D_{\mathrm{max}}=100$, where the maximum normalized error of $\Gamma_{d}$ is shown with respect to the iteration. Since the parameters are very close to the first divergence line, we use a small mixing parameter $\alpha=0.01$, leading to convergence only after 2600 iterations. In accordance with (b), we observe that in the case of $R\leq 6$ the error plateaus at decreasing levels for increasing $R$, which is due to the increased box size reflecting the results from the dense grid calculation. However, for $R\geq 7$ the situation changes, where the leveled off errors are closer together. This means that the error is now governed by the QTT approximation and not by the finite box size anymore. Still, it can be seen that already a maximum bond dimension of 100 is sufficient to reach absolute normalized errors $<10^{-3}$ for larger values of $R$, similar as in (b). Further improving the results can be achieved by increasing the maximum bond dimension.
These calculations were performed on a single 512 GB node on a cluster without parallelization, where in Fig. 12, the runtime of a single iteration of the iterative parquet scheme is shown. For the dense grid calculations performing the same iterative parquet cycle was possible only up to $R=9$ due to the exponentially increasing memory demand. In contrast, using the QTT approach, calculations for $R=11$ were easily carried out on this single node without parallelization. This demonstrates the advantage of QTTs, where memory occupation and operations scale logarithmically with increasing resolution (see Fig. 12(b)), in contrast to the rapid growth in memory and computational costs encountered in standard methods. This allows for efficient computation on large grids, providing a significant advantage over dense grid implementations.
Figure 12: Runtime of a single iteration of the iterative parquet scheme for various maximum bond dimensions and grid size parameters $R$ in the case of $\beta=U=1$. The dashed lines indicate (b) the quartic runtime scaling with $D_{\mathrm{max}}$ (governed by the BSEs), and (c) linear increase with $R$.
@@ VII Conclusion and outlook
This work represents a large step forward in solving many-body problems with quantum field theory methods in QTT representations. The chosen example, the self-consistent solution of parquet equations, is a challenging one, requiring both efficiency in constructing the QTT representation of two-particle vertices and in evaluation of matrix multiplications and variable shifts within this representation. At the same time, the parquet equations for the simplest model, the Hubbard atom, can be solved analytically, allowing for careful benchmarking and assessment of the performance at each step of the solution separately. In this paper we have numerically shown that the QTT representation of the vertex frequency dependence is suitable for solving the parquet equations and that together with TCI it leads to only logarithmic scaling in the grid size and with fourth power in the maximum bond dimension. For the two examples of Hubbard atom and SIAM we observed that the bond dimension of $\sim$100-200 is enough to obtain the solution with high accuracy. For the case of Hubbard atom we see a saturation (or even decrease) of the bond dimension with increasing the inverse temperature $\beta$. We also expect a saturation or only slow growth of the bond dimension with $\beta$ in more general cases, as conjectured in Ref. 31.
Although for explicit testing we have chosen models having no other degrees of freedom than frequencies (no momentum or orbital dependence), the dissection of the parquet equations solver into operations on QTTs—TCI compression and MPO-MPO contractions—is general and the extension to lattice models is straightforward. All results presented here were obtained on a single core with 512 GB memory and the grid sizes in each frequency variable were up to $2^{20}$. This high compressibility of the frequency dependence of vertex functions can in the future be exploited (i) to solve parquet equations for lattice systems with high momentum resolution and orbital degrees of freedom, needed to address material properties; (ii) to apply QTCI to other vertex based methods, such as the functional renormalization group (fRG) [53, 54, 55]; ladder extensions [39] or fRG extensions [56] of dynamical mean-field theory (possibly using as input results for the local vertex obtained using the numerical renormalization group [57, 2, 58]); embedded multi-boson exchange methods [59]; or the Migdal-Eliashberg theory in ab initio calculations [60, 61].
@@ Acknowledgements.
This work was funded in part by the Austrian Science Fund (FWF) Projects No. P 36332 (Grant DOI 10.55776/P36332) and V 1018 (Grant DOI 10.55776/V1018). For open access purposes, the authors have applied a CC BY public copyright license to any author-accepted manuscript version arising from this submission. This work was funded in part by the Deutsche Forschungsgemeinschaft under Germany’s Excellence Strategy EXC-2111 (Project No. 390814868). M.K.R. and J.v.D. acknowledge support from DFG grant LE 3883/2-2. This work is part of the Munich Quantum Valley, supported by the Bavarian state government with funds from the Hightech Agenda Bayern Plus. H.S. was supported by JSPS KAKENHI Grants No. 21H01041, No. 21H01003, and No. 23H03817 as well as JST FOREST Grant No. JPMJFR2232 and JST PRESTO Grant No. JPMJPR2012, Japan. Calculations have been partly performed on the Vienna Scientific Cluster (VSC).
@@ A.2 Bethe–Salpeter picture / Channel native description
Following Ref. 37, in this work we use the following frequency convention for the ph and pp channels
@@ Appendix B Affine transformations
In quantics representation, the tensor $T$ can be factorized to an MPO with small bond dimension, which allows cheap transformation of functions given in QTT format through a single MPO-MPS contraction.
To construct this MPO, it is useful to start with fused rather than interleaved indices, i.e., we only separate out the length scales of the vector, but not its components:
@@ Appendix E SDE in QTT format
In Fig. 16, we show the maximum absolute normalized error of the self energy $\Sigma$ obtained by using QTTs in the SDE. The dashed lines represent the errors of the dense grid calculations, which are due to the finite size of the grid. A qualitatively similar behavior to the BSE can be observed, with the difference that already quite low bond dimensions are sufficient for obtaining very low errors. This is caused by the frequency dependence of the functions in the SDE, where only the full vertex depends on three Matsubara frequencies.
@@
At this point, we should repeat emphasizing the fact that exponentially increasing the number of grid points by increasing the grid parameter $R$ exponentially improves the error, but only comes with linearly increasing computational cost. In the case of $R=14$ the computation with a maximum bond dimension of 100 took only around 100 seconds using QTTs on a single 512 GB node on a cluster, while without the use of QTTs the calculation would include computations with the numerical data of the full vertex, which is the size of $8\times 2^{3\times 14}\simeq 3.5\times 10^{13}$ bytes. This would only be possible by engaging a larger number of nodes on a cluster, which emphasizes the strength of the QTCI approach.

### v3

@@ V.3 Parquet equation
The parquet equation (15) with its frequency shifts as in (16) can be solved entirely in QTT by first converting all the vertex functions $\Lambda,\Phi$ to the required channel as described in the previous section, then performing their summation and subtraction as shown in Ref. [37, Sec. 4.7]. The resultant QTTs for $F$ (and subsequently $\Gamma$ obtained from (13)) are then compressed to a maximum bond dimension $D_{\mathrm{max}}$ in $\mathcal{O}(D_{\mathrm{max}}^{3}R)$ computation time. Hence, the parquet equation has the same $\mathcal{O}(D_{\mathrm{max}}^{3}R)$ computational cost as channel transformations. Further investigation of error and runtime scaling can be found in App. D.
@@ V.4 Bethe–Salpeter equation
where $|_{\omega=\omega^{\prime\prime}}$ denotes the restriction of the result to $\omega=\omega^{\prime\prime}$. Note that Eq. (24a) has the structure of a matrix multiplication in the combined index $(\nu^{\prime},\omega^{\prime})$. Thus, this equation can be evaluated in QTT format through standard MPO-MPO contraction, as is illustrated in Fig. 8(b). For numerical computations, and, thus, also for the QTT approach, the infinite Matsubara sums have to be truncated. The number of Matsubara frequencies taken into account is governed by the chosen Matsubara grid, containing $2^{R}$ points in each direction. Introducing dummy indices has a runtime and memory cost of $\mathcal{O}(D_{\mathrm{max}}^{2}R)$, which is much smaller than the cost of other steps in the algorithm.
@@ V.5 Schwinger–Dyson equation
V.6 Initialization and update strategy
Let us briefly discuss how the iterative parquet calculations are performed in practice. We start by compressing the initial inputs $G_{0}$ and the exact fully irreducible vertices $\Lambda_{r}$ with TCI, where in the Hubbard atom case the exact $\Lambda_{r}$ is taken as input and in the SIAM case we make use of the parquet approximation ($\Lambda_{d}=U,\Lambda_{m}=-U,\Lambda_{s}=2U,\Lambda_{t}=0$). We then set $\Gamma_{r}=\Lambda_{r},F_{d}=U,F_{m}=-U,F_{s}=2U,F_{t}=\Lambda_{t},G=G_{0},\Phi_{r}=0$, where the QTT representation for the input full vertices $F_{r}=U$ can easily be obtained with or without TCI since this can be represented by a QTT of bond dimension one. Then, the right sides of the four BSEs are computed using (SVD-based) MPO-MPO contractions that were discussed in Sec. IV.3. The output of the BSEs is then linearly mixed with the QTTs of the input $\Phi_{r}$ resulting in updated QTT approximations of the reducible vertices $\Phi_{r}$. These updated $\Phi_{r}$ QTTs are then used as input in the parquet equations; as output, the QTT representations of $\Gamma_{r}$ and $F_{r}$ are updated. As a last step the QTT representation of the self-energy is updated by solving the SDE in QTT format. The updated QTT vertices and self-energy are then used as new input in the next iteration step of the iterative parquet scheme.
@@ VI.1 Hubbard atom
The atomic limit of the Hubbard model gives rise to rich two-particle correlations giving insight into strong-copupling limit of the Hubbard model [1, 51, 40]. This atomic limit offers considerable simplification, as the vertex functions are analytically known and become independent of momentum [40]. Therefore, the Hubbard atom serves as an ideal first test case for exploring two-particle properties in strongly correlated electron systems using the QTT framework.
@@
Figure 11: (a) Maximum absolute normalized error $||\Delta\Gamma_{d}||_{\infty}:=||\Gamma_{d,\text{iterative-parquet}}-\Gamma_{d,\text{exact}}||_{\infty}/||\Gamma_{d,\text{exact}}||_{\infty}$ of the iterative parquet scheme with QTCI for the irreducible vertex $\Gamma_{d}$ compared to the exact values for various grid sizes ($2^{3R}$ grid points) in the case of $U=\beta=1$, plotted (a) as a function of $D_{\mathrm{max}}$ after 30 iterations, and as a function of iteration number for (b) $D_{\mathrm{max}}=100$ and (c) $D_{\mathrm{max}}=200$. Dashed lines indicate the results from the dense grid calculation without QTCI. (d) The maximum absolute normalized error of $\Gamma_{d}$, shown close to the first divergence, for $\beta=1.55,U=2.3$ and $D_{\mathrm{max}}=100$, up to a very large number of 2600 iterations.
@@
Next, we show that computations can also be performed for a more challenging case with the parameters $\beta=1.55,U=2.3$, which are chosen in the same way as in Ref. [13]. This case is interesting since $\beta U=3.565$ is very close to the point, where the irreducible vertex in the density channel diverges ($\beta U\approx 3.628$[45, 56, 57, 58, 59, 60]. We show the results of these calculations in Fig. 11(d) for a maximum bond dimension $D_{\mathrm{max}}=100$, where the maximum normalized error of $\Gamma_{d}$ is shown with respect to the iteration. Since the parameters are very close to the first divergence line, we use a small mixing parameter $\alpha=0.01$, leading to convergence only after 2600 iterations. In accordance with (b), we observe that in the case of $R\leq 6$ the error plateaus at decreasing levels for increasing $R$, which is due to the increased box size reflecting the results from the dense grid calculation. However, for $R\geq 7$ the situation changes, where the leveled off errors are closer together. This means that the error is now governed by the QTT approximation (i.e. by $D_{\mathrm{max}}$) and not by the finite box size anymore. Still, it can be seen that already a maximum bond dimension of 100 is sufficient to reach absolute normalized errors $<10^{-3}$ for larger values of $R$, similar as in (b).
The above analysis demonstrates that during BSE iterations, errors due to the QTT representation do not accumulate – if they did, the solid lines in Figs. 11(b,c) would slope upwards with increasing iteration number. Instead, the error saturates, at a value governed by the maximum of the error due to the Matsubara sum truncation and the initial QTT approximation. The accuracy of the results can be improved systematically by increasing the maximum bond dimension. In principle, this can also be done during the course of the BSE iterations if these generate structures of increasing complexity, but that was not necessary for the calculations presented here.
Our calculations were performed on a single 512 GB node on a cluster without parallelization, where in Fig. 12, the runtime of a single iteration of the iterative parquet scheme is shown. For the dense grid calculations performing the same iterative parquet cycle was possible only up to $R=9$ due to the exponentially increasing memory demand. In contrast, using the QTT approach, calculations for $R=11$ were easily carried out on this single node without parallelization. This demonstrates the advantage of QTTs, where memory occupation and operations scale logarithmically with increasing resolution (see Fig. 12(b)), in contrast to the rapid growth in memory and computational costs encountered in standard methods. This allows for efficient computation on large grids, providing a significant advantage over dense grid implementations.
Figure 12: Runtime of a single iteration of the iterative parquet scheme for various maximum bond dimensions and grid size parameters $R$ in the case of $\beta=U=1$. The dashed lines indicate (a) the quartic runtime scaling with $D_{\mathrm{max}}$ (governed by the BSEs), and (b) linear increase with $R$.
@@ VII Conclusion and outlook
This work represents a large step forward in solving many-body problems with quantum field theory methods in QTT representations. The chosen example, the self-consistent solution of parquet equations, is a challenging one, requiring both efficiency in constructing the QTT representation of two-particle vertices and in evaluation of matrix multiplications and variable shifts within this representation. At the same time, the parquet equations for the simplest model, the Hubbard atom, can be solved analytically, allowing for careful benchmarking and assessment of the performance at each step of the solution separately. In this paper we have numerically shown that the QTT representation of the vertex frequency dependence is suitable for solving the parquet equations and that together with TCI it leads to only logarithmic scaling in the grid size and with fourth power in the maximum bond dimension. For the two examples of Hubbard atom and SIAM we observed that the bond dimension of $\sim$100-200 is enough to obtain the solution with high accuracy. For the case of Hubbard atom we see a saturation (or even decrease) of the bond dimension with increasing the inverse temperature $\beta$. We also expect a saturation or only slow growth of the bond dimension with $\beta$ in more general cases, as conjectured in Ref. 34.
The naive iteration of the parquet equations is difficult to converge in some regimes, cf. Sec. III.6. While this problem is almost orthogonal to questions of representation and compression of the vertices, QTTs offer potential synergies: in particular, the greatly reduced size of QTT vertices may enable a solver to keep a “convergence history” and use non-linear mixing schemes, which have been shown to stabilize convergence in Hartree-Fock [61], or quasi-Newton solvers, which have been able to access previously hidden solutions in self-consistent diagrammatic theories [62].
Although for explicit testing we have chosen models having no other degrees of freedom than frequencies (no momentum or orbital dependence), the dissection of the parquet equations solver into operations on QTTs—TCI compression and MPO-MPO contractions—is general and the extension to lattice models is straightforward. All results presented here were obtained on a single core with 512 GB memory and the grid sizes in each frequency variable were up to $2^{20}$. This high compressibility of the frequency dependence of vertex functions can in the future be exploited (i) to solve parquet equations for lattice systems with high momentum resolution and orbital degrees of freedom, needed to address material properties (see App. G for a brief discussion of how to deal with such additional degrees of freedom); (ii) to apply QTCI to other vertex based methods, such as the functional renormalization group (fRG) [63, 64, 65]; ladder extensions [42] or fRG extensions [66] of dynamical mean-field theory (possibly using as input results for the local vertex obtained using the numerical renormalization group [67, 2, 68]); embedded multi-boson exchange methods [69]; or the Migdal-Eliashberg theory in ab initio calculations [70, 71].
@@ Acknowledgements.
This work was funded in part by the Austrian Science Fund (FWF) Projects No. P 36332 (Grant DOI 10.55776/P36332) and V 1018 (Grant DOI 10.55776/V1018). For open access purposes, the authors have applied a CC BY public copyright license to any author-accepted manuscript version arising from this submission. This work was funded in part by the Deutsche Forschungsgemeinschaft under Germany’s Excellence Strategy EXC-2111 (Project No. 390814868). M.K.R. and J.v.D. acknowledge support from DFG grant LE 3883/2-2. This work is part of the Munich Quantum Valley, supported by the Bavarian state government with funds from the Hightech Agenda Bayern Plus. H.S. was supported by JSPS KAKENHI Grants No. 21H01041, No. 21H01003, No. 22KK0226, and No. 23H03817 as well as JST FOREST Grant No. JPMJFR2232 and JST PRESTO Grant No. JPMJPR2012, Japan. Calculations have been partly performed on the Vienna Scientific Cluster (VSC) using the Quantics.jl and TensorCrossInterpolation.jl libraries. The raw data for the figures reported are available at Ref. 74.
@@ A.2 Bethe–Salpeter picture / Channel native description
Following Ref. 40, in this work we use the following frequency convention for the ph and pp channels
@@ Appendix B Affine transformations
In quantics representation, the tensor $T$ can be factorized to an MPO with small bond dimension, which allows cheap transformation of functions given in QTT format through a single MPO-MPS contraction. To construct this MPO, it is useful to start with fused rather than interleaved indices, i.e., we only separate out the length scales of the vector, but not its components:
@@ Appendix E SDE in QTT format
In Fig. 16, we show the maximum absolute normalized error of the self-energy $\Sigma$ obtained by using QTTs in the SDE. The dashed lines represent the errors of the dense grid calculations, which are due to the finite size of the grid. A qualitatively similar behavior to the BSE can be observed, with the difference that already quite low bond dimensions are sufficient for obtaining very low errors. This is caused by the frequency dependence of the functions in the SDE, where only the full vertex depends on three Matsubara frequencies.
@@
At this point, we should repeat emphasizing the fact that exponentially increasing the number of grid points by increasing the grid parameter $R$ exponentially reduces the error (exponential convergence to the true solution), but only comes with linearly increasing computational cost. In the case of $R=14$ the computation with a maximum bond dimension of 100 took only around 100 seconds using QTTs on a single 512 GB node on a cluster, while without the use of QTTs the calculation would include computations with the numerical data of the full vertex, which is the size of $8\times 2^{3\times 14}\simeq 3.5\times 10^{13}$ bytes. This would only be possible by engaging a larger number of nodes on a cluster, which emphasizes the strength of the QTCI approach.
@@ Appendix F Technical limitations
Appendix G Model extensions
In future work, our goal will be to apply the QTT representation to two-particle calculations with orbital and momentum degrees of freedom. Suppose that these are labeled by an additional (composite) index, say $i=1,\dots,N$, then the vertex carries four such indices, $F_{ijkl}$, and has $N^{4}$ components. The memory costs for such computations depend on how QTTs are used to parametrize the vertex.
For example, a naive approach would be to use a separate QTT to parameterize the frequency dependence of $F_{ijkl}(\nu,\nu^{\prime},\omega)$ for each index combination $(i,j,k,l)$. This would require $N^{4}$ different QTTs and $N^{4}$ BSEs connecting them all, etc. We estimate that with this approach, computations for $N=6$, $R=10$ and $D_{\mathrm{max}}=200$ should be feasible on a single 512 GB node.
However, such a naive approach would not exploit low-rank structures that may arise if different vertex components have similar frequency dependencies. In such a case, it could be more efficient to use a single QTT to parametrize the dependence of the vertex on its frequencies and all $i$-indices. To pursue such a strategy and optimize its efficiency, further methodological developments will be required to address some open questions: What is the best grouping and order of quantics indices for a combined frequency and momentum/orbital representation? How can MPO-MPO contractions (the current bottleneck) be performed more efficiently? What are the best strategies for parallelizing the computations? These issues are currently being explored in ongoing work.